# What sets the metallicity of Ultra-Faint Dwarfs?

### v1

@@ 3.1.3 Effect of the accreted IGM metallicity on the average stellar metallicity in UFDs
To this end, we sample the $Z_{\rm IGM}$ distributions constructed in the previous section, interpolated between redshifts, for every redshift step of the MAH $({\sim}100\ {\rm Myr})$ for each GRUMPY model galaxy. The sampled values are taken as the metallicity of the accreted gas instead of assuming a constant value as in the fiducial model calculation. Therefore, the metallicity of accreted IGM varies with redshift, and from galaxy to galaxy. Figure 3 shows the resulting average stellar metallicities as a function of galaxy $V$-band absolute magnitude. In this section and 3.2, we define model metallicity by $Z_{\star}$ interchangeably with $\mathrm{[Fe/H]}$ using the GRUMPY base model as in Kravtsov & Manwadkar (2022); Manwadkar & Kravtsov (2022), noting that the model instantaneously recycles stellar yields. This choice does not effect our conclusions, as at early times [Fe/H] would be lower than Z because of the low contribution of SNIa, strengthening our results. The metallicity of the accreted gas has little effect for bright ($M_{V}<-7$) galaxies, but affects the average metallicity of UFD galaxies substantially. In particular, models with the IGM metallicity sampled from the simulation distributions produce metallicities considerably lower than the fiducial model with constant $Z_{\rm IGM}=10^{-3}\,Z_{\odot}$ and lower than the typical metallicities of observed UFDs. The relations of these models can be matched well by the model with constant $Z_{\rm IGM}=10^{-4}\,Z_{\odot}$. In fact, the predicted ${\rm[Fe/H]}-M_{V}$ relation is insensitive to the IGM metallicity or its distribution for $Z_{\rm IGM}\lesssim 10^{-4}\,Z_{\odot}$.
@@
Figure 3: Average stellar metallicity vs V-band magnitude for observed and model galaxies in this study. Lines are the median, and shaded or hatched regions contain $68\%$ of model galaxies for a given $M_{V}$. Red points are observed MW dwarf galaxies reported in Pace (2024) which have mean spectroscopic metallicities derived using $>5$ stars. All model parameters other than $Z_{\rm IGM}$ are the fiducial values of Manwadkar & Kravtsov (2022). For models informed by cosmological simulations, shown in violet, $Z_{\rm IGM}$ is drawn from the distribution categorized in Section 3.1.2. The MK22 and $Z_{\rm IGM}$$=-4$ models use a constant value. All simulation-informed models, which have very little enriched gas available to UFDs, possess lower average metallicity than observed UFDs, and the intrinsic scatter in all models is significantly less than the intrinsic scatter of observed galaxies. $Z_{\rm IGM}$$=-4$ is included to demonstrate the behavior of a typical metallicity floor, and all simulation-informed models very closely follow this behavior in both the average and scatter regardless of metallicity floor or prescription.
@@
Figure 4: Metallicity vs $V$-band absolute magnitude relation of galaxies in the models that use S24 and FIRE-2 IGM metallicity distributions and in the model with constant $Z_{\rm IGM}=10^{-4}\,Z_{\odot}$. Lines show the medians of model galaxies in bins of absolute magnitude and shaded bands contain $95\%$ of galaxies in a particular model; galaxies outside the 95% bands are shown as individual dots. This figure shows that the luminosity–metallicity relation is generally insensitive to the form of $Z_{\rm IGM}$ distribution for $Z_{\rm IGM}<10^{-4}\,Z_{\odot}$. However, the simulation-informed models produce a tail of galaxies enriched considerably more than the median for their luminosity, which is lacking in the constant $Z_{\rm IGM}=10^{-4}\,Z_{\odot}$ model. Likewise, the S24-informed model produces a larger fraction of galaxies in the high-metallicity tail than the FIRE-2 informed model because $Z_{\rm IGM}$ in the former is wider and extends to higher metallicities (see Figure 2).
@@
Finally, the models predict a much smaller scatter about the median ${\rm[Fe/H]}-M_{V}$ relation than exhibited by observed UFD galaxies. Figure 4 shows that although the luminosity–metallicity relation for model galaxies with $Z_{\rm IGM}$ sampled from cosmological simulations is consistent with the constant $Z_{\rm IGM}=10^{-4}\,Z_{\odot}$ models, drawing $Z_{\rm IGM}$ from these extended distributions does produce differences in the tails. Simulation-informed $Z_{\rm IGM}$ models can produce a galaxy population with a wide range in metallicities, however the number of such galaxies is small enough that it contributes to a far reaching but weak tail of an otherwise very narrowly peaked distribution inconsistent with the observed scatter in UFD metallicities.
@@
The tail for S24 $Z_{\rm IGM}$ models is significantly stronger than that of FIRE-2 m12 $Z_{\rm IGM}$ models, due to the greater contribution of $\mathrm{[Fe/H]}\gtrsim-2$ gas available to UFDs in the S24 simulation. By selectively choosing which redshifts to sample from this distribution, it was found that the metallicity of fainter galaxies near the plateau are almost entirely dictated by the metallicity distribution at redshifts above the fiducial reionization redshift $z\approx 8.5$, while the extent of brighter galaxy metallicity was similar to that of FIRE-2 m12 $Z_{\rm IGM}$ models. Therefore, it is the extremely broad S24 $Z_{\rm IGM}$ distribution at lower redshifts which contributes to the wide tail for $M_{V}<-5$ galaxies shown in Figure 4. The metallicity distributions in MEGATRON simulations are narrower than in S24 but produce similar metallicity scatter as described for sampling only the high redshift S24 $Z_{\rm IGM}$ distribution. However, MEGATRON simulations at lower redshifts are not yet available to compare over the full redshift range.
@@
Note that although the constant $Z_{\rm IGM}=10^{-3}\,Z_{\odot}$ model can match the average metallicity of the faint observed UFDs, such a model does not reproduce the observed scatter. Moreover, such a high metallicity floor of the accreted gas is excluded by the observation of UFD stars with metallicities further than one dex below this value (e.g., Simon, 2019). Although a wide distribution with a central value near $\mathrm{[Fe/H]}=-3$ would be consistent with observations, such a distribution does not seem to be produced by modern state-of-the-art galaxy formation simulations. Therefore, we conclude that variability in the pre-enrichment of the IGM gas accreted by UFDs cannot explain the metallicity of these systems.
@@
This conclusion is consistent with the results of modern simulations of dwarf galaxies in the cosmological context, which underpredict metallicities of UFDs by $\sim 0.5-1$ dex (e.g., see Fig. 1 in Sanati et al., 2023, or Fig. 6 in Go et al. 2025). This means that the gas accreted by UFD progenitors in these simulations is not enriched to the level that would result in a metallicity floor which would create a plateau at the average UFD stellar metallicity $\mathrm{[Fe/H]}\approx-2.5$. As we show next, this also indicates that the feedback-driven outflows in UFD progenitors in simulations are likely too strong.
@@ 3.2 Effect of mass loading factor on the stellar metallicity of UFDs
The relatively tight bright-end luminosity–metallicity relation constrains most of the parameters of equation 3. As shown in Figure 3, the model reproduces the relation for galaxies of $M_{V}\lesssim-6$ well using these parameters. The metallicity of UFD galaxies at lower luminosities is approximately constant on average and such average behavior can be matched if the average metallicity of the accreted gas was set to $Z_{\rm IGM}\approx 10^{-3}\,Z_{\odot}$ or greater. As was shown above, however, the accreted gas metallicity range predicted by several state-of-the-art simulations is an order of magnitude lower. In this case, the average metallicity of UFD galaxies can be matched by decreasing the $\eta_{\rm max}$  value, which results in retention of a larger fraction of the metals which model UFDs produce within their ISM.
@@
Figure 5: Same as Figure 3, however the GRUMPY model galaxies shown now vary $\eta_{\rm max}$ with fixed $Z_{\rm IGM}=10^{-4}\,Z_{\odot}$, a value which produces outputs consistent with the $Z_{\rm IGM}$ distributions categorized in all cosmological simulation-informed models studied in previous sections. Varying $\eta_{\rm max}$ has a profound effect on the location of the plateau, and $2000200<\eta_{\rm max}<$ 2000 is sufficient to bracket the entire observed range in average metallicity. Reducing $\eta_{\rm max}$ from the fiducial value of $2000$ to $1000$ almost exactly recovers the same plateau position as the $Z_{\rm IGM}=10^{-3}\,Z_{\odot}$ model, and the plateau of the $\eta_{\rm max}=650$ model is well aligned with the median observed UFD metallicity of $\mathrm{[Fe/H]}=-2.48$.
@@ 3.3 Comparison of model and observed stellar metallicity distributions in UFDs
Figure 6: Comparing the shape of the MDF for models with various $Z_{\rm IGM}$ and $\eta_{\rm max}$ prescriptions to observed UFD MDFs. Each dashed line in the left 3 columns is the MDF for an individual galaxy which falls within $1\sigma$ of the magnitude of the observed galaxy to which it is compared. Here we compare only the shape of the MDF by offsetting the median value of the model MDF to match that of observation. Each stellar population of a model MDF is convolved with a stellar metallicity error from the observed galaxy in a bootstrap process. Each row shows a different observed UFD with MDF as a thick red line, while each column shows a different model. The number of spectroscopic observed stars making the MDF is given below the name of the UFD. The right column shows the cumulative distribution of Kolmogorov-Smirnov probability values comparing each model to the observed galaxy within a row. All model prescriptions produce MDF shapes that are nearly identical in comparison to observation, therefore no model is strongly preferred or disfavored. The $\eta_{\rm max}$$=200$ model has greater scatter and stronger tails than other distributions, as expected from the width of the contour in Figure 5. Only in the faintest galaxies does this model stand out in KS probability, due to the weakly sampled tail of the observed MDFs in these galaxies.
@@
[…]
@@
Comparing to the $Z_{\rm IGM}$$=-4$, $\eta_{\rm max}=2000$ model, it can be seen that both other models shown have a greater variety of individual MDF shapes. The $\eta_{\rm max}=200$ model produces more MDFs with significantly stronger tails, especially at the lowest metallicities. This is consistent with the wider $1\sigma$ contour in luminosity–metallicity for this model shown in Figure 5. The model with the S24-based $Z_{\rm IGM}$ distribution, which produces the strongest tail of high metallicities in galaxies of a given $M_{V}$ (see Figure 4) does not produce notably more MDFs skewed to high metallicity than other models. Overall, there is no clear qualitative difference in the average MDF between models, or their success in matching the observed MDFs.
@@ 4.1 The role of Pop III stars in chemical enrichment of UFD stars
Much of the current discussion of the pre-enrichment of the IGM in the literature is centered on the physics of very low metallicity and Pop III stars. In our study, the FIRE-2 and S24 simulations do not model Pop III stars and their contribution to feedback and enrichment, while the MEGATRON models do include explicit modeling of Pop III stars and their contribution to chemical enrichment. We find that there is very little difference in the luminosity–metallicity relation of dwarf galaxies when accretable IGM metallicity distributions from these simulations are used in the GRUMPY galaxy formation model. Likewise, there is no significant difference in the predicted metallicity distributions of stars resulting from these distributions. The exact chemical abundance of stars may differ due to Pop III yields, however we do not model the accretion of individual elements in detail.
At redshifts $z\gtrsim 8$, where UFDs are most likely to form the bulk of their stars, the distributions of IGM available to UFDs are qualitatively similar, in that very little gas is enriched above $\mathrm{[Fe/H]}=-4$. The contribution of Pop III stars explicitly modeled in the MEGATRON simulations thus does not have substantial amounts of more enriched gas available compared to the FIRE-2 and S24 simulations. This is consistent with observations of the lowest metallicity DLA system at $z=3.08$, which has an upper limit on metallicity of $\mathrm{[Fe/H]}\lesssim-3.66$ and constraints on the abundances of C, Al, and Si consistent with the origin in a single Pop III supernova (Welsh et al., 2023, see also Erni et al. 2006; Cooke et al. 2011).
@@
Conversely, Prgomet et al. (2022) find that introducing a metallicity dependent IMF to a simulation of a single dwarf galaxy from the EDGE simulations is capable of producing UFDs with stellar metallicities $\mathrm{[Fe/H]}>-2.5$ in agreement with the observed luminosity–metallicity relation. A greater number of high mass stars increases burstiness, regulating star formation and reducing stellar mass for a given halo mass compared to the fiducial model. More high mass stars also increase iron yields commensurately to produce a relatively flat trend for permutations of this galaxy. A similar result was found with the inclusion of radiative feedback in addition to supernovae.

### v2

@@ 3.1.3 Effect of the accreted IGM metallicity on the average stellar metallicity in UFDs
To this end, we sample the $Z_{\rm IGM}$ distributions constructed in the previous section, interpolated between redshifts, for every point in the mass assembly history which GRUMPY samples $(\Delta a\ {\sim}\ 0.014\ )$ for each GRUMPY model galaxy. The sampled values are taken as the metallicity of the accreted gas for that period, instead of assuming a constant value as in the fiducial model calculation. Therefore, the metallicity of accreted IGM varies with redshift, and from galaxy to galaxy.
@@
Figure 3 shows the resulting average stellar metallicities as a function of galaxy $V$-band absolute magnitude. In this section and 3.2, we define model metallicity by $Z_{\star}$ interchangeably with $\mathrm{[Fe/H]}$ using the GRUMPY base model as in Kravtsov & Manwadkar (2022); Manwadkar & Kravtsov (2022), noting that the model instantaneously recycles stellar yields. This choice does not effect our conclusions, as at early times [Fe/H] would be lower than Z because of the low contribution of SNIa, strengthening our results. The metallicity of the accreted gas has little effect for bright ($M_{V}<-7$) galaxies, but affects the average metallicity of UFD galaxies substantially. In particular, models with the IGM metallicity sampled from the simulation distributions produce metallicities considerably lower than the fiducial model with constant $Z_{\rm IGM}=10^{-3}\,Z_{\odot}$ and lower than the typical metallicities of observed UFDs. The relations of these models can be matched well by the model with constant $Z_{\rm IGM}=10^{-4}\,Z_{\odot}$. In fact, the predicted ${\rm[Fe/H]}-M_{V}$ relation is insensitive to the IGM metallicity or its distribution for $Z_{\rm IGM}\lesssim 10^{-4}\,Z_{\odot}$.
@@
Figure 3: Average stellar metallicity vs V-band magnitude for observed and model galaxies in this study. Lines are the median, and shaded or hatched regions contain $68\%$ of model galaxies for a given $M_{V}$. Red points are observed MW dwarf galaxies reported in Pace (2024) which have mean spectroscopic metallicities derived using >5>5 stars. All model parameters other than $Z_{\rm IGM}$ are the fiducial values of Manwadkar & Kravtsov (2022). For models informed by cosmological simulations, shown in violet, $Z_{\rm IGM}$ is drawn from the distribution categorized in Section 3.1.2. The MK22 and $Z_{\rm IGM}$=4=-4 models use a constant value. All simulation-informed models, which have very little enriched gas available to UFDs, possess lower average metallicity than observed UFDs, and the intrinsic scatter in all models is significantly less than the intrinsic scatter of observed galaxies. $Z_{\rm IGM}$=4=-4 is included to demonstrate the behavior of a typical metallicity floor, and all simulation-informed models very closely follow this behavior in both the average and scatter regardless of metallicity floor or prescription.
@@
Figure 4: Metallicity vs $V$-band absolute magnitude relation of galaxies in the models that use S25 and FIRE-2 IGM metallicity distributions and in the model with constant $Z_{\rm IGM}=10^{-4}\,Z_{\odot}$. Lines show the medians of model galaxies in bins of absolute magnitude and shaded bands contain $95\%$ of galaxies in a particular model; galaxies outside the 95% bands are shown as individual dots. This figure shows that the luminosity–metallicity relation is generally insensitive to the form of $Z_{\rm IGM}$ distribution for $Z_{\rm IGM}<10^{-4}\,Z_{\odot}$. However, the simulation-informed models produce a tail of galaxies enriched considerably more than the median for their luminosity, which is lacking in the constant $Z_{\rm IGM}=10^{-4}\,Z_{\odot}$ model. Likewise, the S25-informed model produces a larger fraction of galaxies in the high-metallicity tail than the FIRE-2 informed model because $Z_{\rm IGM}$ in the former is wider and extends to higher metallicities (see Figure 2).
@@
Finally, the models predict a much smaller scatter about the median ${\rm[Fe/H]}-M_{V}$ relation than exhibited by observed UFD galaxies. Figure 4 shows that although the luminosity–metallicity relation for model galaxies with $Z_{\rm IGM}$ sampled from cosmological simulations is consistent with the constant $Z_{\rm IGM}=10^{-4}\,Z_{\odot}$ models, drawing $Z_{\rm IGM}$ from these extended distributions does produce differences in the tails. Simulation-informed $Z_{\rm IGM}$ models can produce a galaxy population with a wide range in metallicities, but the number of such galaxies is small enough that it contributes to a far-reaching but weak tail of an otherwise very narrowly peaked distribution inconsistent with the observed scatter in UFD metallicities.
@@
The tail for S25 $Z_{\rm IGM}$ models is significantly stronger than that of FIRE-2 m12 $Z_{\rm IGM}$ models, due to the greater contribution of $\mathrm{[Fe/H]}\gtrsim-2$ gas available to UFDs in the S25 simulation. By selectively choosing which redshifts to sample from this distribution, it was found that the metallicity of fainter galaxies near the plateau are almost entirely dictated by the metallicity distribution at redshifts above the fiducial reionization redshift $z\approx 8.5$, while the extent of brighter galaxy metallicity was similar to that of FIRE-2 m12 $Z_{\rm IGM}$ models. Therefore, it is the extremely broad S25 $Z_{\rm IGM}$ distribution at lower redshifts which contributes to the wide tail for $M_{V}<-5$ galaxies shown in Figure 4. The metallicity distributions in MEGATRON simulations are narrower than in S25, but produce similar metallicity scatter as described for sampling only the high redshift S25 $Z_{\rm IGM}$ distribution. However, MEGATRON simulations at lower redshifts are not yet available to compare over the full redshift range.
Note that the tail of galaxies with metallicities smaller than the median is generally small in all of the simulations. This is not due to the characteristic metallicity of the accreted IGM, which is far lower than the metallicities in this tail in most of the considered models. Rather, the tail is shaped by the distribution of star formation, and hence outflow, histories of the model galaxies, which reflect differences in the mass assembly histories of their halos.
@@
Although the constant $Z_{\rm IGM}=10^{-3}\,Z_{\odot}$ model can match the average metallicity of the faint observed UFDs, such a model does not reproduce the observed scatter. Moreover, such a high metallicity floor of the accreted gas is excluded by the observation of UFD stars with metallicities further than one dex below this value (e.g., Simon, 2019). Although a wide distribution with a central value near $\mathrm{[Fe/H]}=-3$ would be consistent with observations, such a distribution does not seem to be produced by modern state-of-the-art galaxy formation simulations.
@@
Overall, we conclude that variability in the pre-enrichment of the IGM gas accreted by UFDs cannot explain the metallicity of these systems. This conclusion is consistent with the results of modern simulations of dwarf galaxies in the cosmological context, which underpredict metallicities of UFDs by $\sim 0.5-1$ dex (e.g., see Fig. 1 in Sanati et al., 2023, or Fig. 6 in Go et al. 2025). This means that the gas accreted by UFD progenitors in these simulations is not enriched to the level that would result in a metallicity floor which would create a plateau at the average UFD stellar metallicity $\mathrm{[Fe/H]}\approx-2.5$. As we show next, this also indicates that the feedback-driven outflows in UFD progenitors in simulations are likely too strong.
@@ 3.2 Effect of mass loading factor on the stellar metallicity of UFDs
The relatively tight bright-end luminosity–metallicity relation constrains most of the parameters of equation 3. As shown in Figure 3, the model reproduces the relation for galaxies of $M_{V}\lesssim-6$ well using these parameters. The metallicity of UFD galaxies at lower luminosities is approximately constant on average and such average behavior can be matched if the average metallicity of the accreted gas was set to $Z_{\rm IGM}\approx 10^{-3}\,Z_{\odot}$ or greater. As was shown above, however, the accreted gas metallicity range predicted by several state-of-the-art simulations is consistent with average an order of magnitude lower. In this case, the average metallicity of UFD galaxies can be matched by decreasing the $\eta_{\rm max}$  value, which results in retention of a larger fraction of the metals which model UFDs produce within their ISM.
@@
Figure 5: Same as Figure 3, however the GRUMPY model galaxies shown now vary $\eta_{\rm max}$ with fixed $Z_{\rm IGM}=10^{-4}\,Z_{\odot}$, a value which produces outputs consistent with the $Z_{\rm IGM}$ distributions categorized in all cosmological simulation-informed models studied in previous sections. Varying $\eta_{\rm max}$ has a profound effect on the location of the plateau, and $200<\eta_{\rm max}<2000$ is sufficient to bracket the entire observed range in average metallicity. Reducing $\eta_{\rm max}$ from the fiducial value of $2000$ to $1000$ almost exactly recovers the same plateau position as the $Z_{\rm IGM}=10^{-3}\,Z_{\odot}$ model, and the plateau of the $\eta_{\rm max}=650$ model is well aligned with the median observed UFD metallicity of $\mathrm{[Fe/H]}=-2.48$.
@@ 3.3 Comparison of model and observed stellar metallicity distributions in UFDs
Figure 6: Comparing the shape of the MDF for models with various $Z_{\rm IGM}$ and $\eta_{\rm max}$ prescriptions to observed UFD MDFs. Each dashed line in the left 3 columns is the MDF for an individual galaxy which falls within $1\sigma$ of the magnitude of the observed galaxy to which it is compared. Here we compare only the shape of the MDF by offsetting the median value of the model MDF to match that of observation. Each stellar population of a model MDF is convolved with a stellar metallicity error from the observed galaxy in a bootstrap process. Each row shows a different observed UFD with MDF as a thick red line, while each column shows a different model. The number of spectroscopic observed stars making the MDF is given below the name of the UFD. The right column shows the cumulative distribution of Kolmogorov-Smirnov probability values comparing each model to the observed galaxy within a row. All model prescriptions produce MDF shapes that are nearly identical in comparison to observation, therefore no model is strongly preferred or disfavored. The $\eta_{\rm max}$=200=200 model has greater scatter and stronger tails than other distributions, as expected from the width of the contour in Figure 5. Only in the faintest galaxies does this model stand out in KS probability, due to the weakly sampled tail of the observed MDFs in these galaxies.
@@
To quantify the relative success of different models in reproducing the observed MDF shape, we perform a Kolmogorov-Smirnov test comparing observed MDFs to the individual galaxy MDFs for each model. The resulting distribution of KS probability values for galaxies within a model gives a quantitative comparison of which model better reproduces the observed MDF shape. We compare models to $5$ observed UFD MDFs, which have metallicity measurements of at least $20$ stars and do not have significant overlap in $M_{V}$. Figure 6 shows MDFs from models with $Z_{\rm IGM}=10^{4}\,Z_{\odot}$, $\eta_{\rm max}=200\ {\rm and}\ 2000$, and a model with $Z_{\rm IGM}$ drawn from the S25-based $Z_{\rm IGM}$ distribution with $\eta_{\rm max}=650$. The first two span the range in $\eta_{\rm max}$ required to bracket the observed luminosity–metallicity relation. The latter gives an intermediate value of $\eta_{\rm max}$ which creates a luminosity–metallicity curve approximating the mean metallicity of UFDs and shows a non constant, simulation-informed $Z_{\rm IGM}$.
@@
Comparing to the $Z_{\rm IGM}$=4=-4, $\eta_{\rm max}=2000$ model, it can be seen that both other models shown have a greater variety of individual MDF shapes. The $\eta_{\rm max}=200$ model produces more MDFs with significantly stronger tails, especially at the lowest metallicities. This is consistent with the wider $1\sigma$ contour in luminosity–metallicity for this model shown in Figure 5. The model with the S25-based $Z_{\rm IGM}$ distribution, which produces the strongest tail of high metallicities in galaxies of a given $M_{V}$ (see Figure 4) does not produce notably more MDFs skewed to high metallicity than other models. Overall, there is no clear qualitative difference in the average MDF between models, or their success in matching the observed MDFs.
@@ 4.1 The role of Pop III stars in chemical enrichment of UFD stars
Much of the current discussion of the pre-enrichment of the IGM in the literature is centered on the physics of very low metallicity and Pop III stars. In our study, the FIRE-2 and S25 simulations do not model Pop III stars and their contribution to feedback and enrichment, while the MEGATRON models do include explicit modeling of Pop III stars and their contribution to chemical enrichment. We find that there is very little difference in the luminosity–metallicity relation of dwarf galaxies when accretable IGM metallicity distributions from these simulations are used in the GRUMPY galaxy formation model. Likewise, there is no significant difference in the predicted metallicity distributions of stars resulting from these distributions. The exact chemical abundance of stars may differ due to Pop III yields, however we do not model the accretion of individual elements in detail.
At redshifts $z\gtrsim 8$, where UFDs are most likely to form the bulk of their stars, the distributions of IGM available to UFDs are qualitatively similar, in that very little gas is enriched above $\mathrm{[Fe/H]}=-4$. The contribution of Pop III stars explicitly modeled in the MEGATRON simulations thus does not lead to substantial amounts of more enriched gas available compared to the FIRE-2 and S25 simulations. This is consistent with observations of the lowest metallicity damped Lyman $\alpha$ system at $z=3.08$, which has an upper limit on metallicity of $\mathrm{[Fe/H]}\lesssim-3.66$ and constraints on the abundances of C, Al, and Si consistent with the origin in a single Pop III supernova (Welsh et al., 2023, see also Erni et al. 2006; Cooke et al. 2011).
@@
Conversely, Prgomet et al. (2022) find that introducing a metallicity dependent IMF to a simulation of a single dwarf galaxy from the EDGE simulations is capable of producing UFDs with stellar metallicities $\mathrm{[Fe/H]}>-2.5$ in agreement with the observed luminosity–metallicity relation. A greater number of high mass stars increases burstiness, regulating star formation and reducing stellar mass for a given halo mass compared to their fiducial model. More high mass stars also increase iron yields commensurately to produce a relatively flat trend for permutations of this galaxy. A similar result was found with the inclusion of radiative feedback in addition to supernovae.